OpenStudy (anonymous):

What is the PH of a solution whose H3O+ concentration is 10.0 m? Is this solutoin acidic or basic?

OpenStudy (misssunshinexxoxo):

1.0 * 10-4 is the OH- concentration [H30+] = 10(-pH) = 10-10 = 1.0 * 10-10 M [OH-] = (Kw) / [H30+] = (1.0 * 10-14) /(1.0 * 10-10) = 1.0 * 10-4 M

OpenStudy (misssunshinexxoxo):

Always acidic

OpenStudy (anonymous):

Thank you so much

OpenStudy (misssunshinexxoxo):

Please mark me best response:)

OpenStudy (misssunshinexxoxo):

for a medal Welcome to openstudy